OpenStudy (anonymous):

i need help! i am supposed to simplify the cube root of 9 divided by the 5th root of 27

OpenStudy (anonymous):

³√ 9 = ³√ 3² ⁵√27 = ⁵√ 3³ @dawnd4fun Can you divide them now?

OpenStudy (anonymous):

i know the answer is 3to the root of 15 i can not see how to get there.

OpenStudy (anonymous):

In the exponential form: = 3^ (2/3) - 3 ^(3/5) = ....

OpenStudy (anonymous):

I am just too stupid

OpenStudy (anonymous):

Since you've already had the same base 3 just subtract 2/3 - 3/5 = ..

OpenStudy (anonymous):

how does that equal 3 to the root 15

OpenStudy (anonymous):

the answer is in the book

OpenStudy (anonymous):

10/15 minus 9/15 = 1/15

OpenStudy (anonymous):

There you go, now switch back to the root form!

OpenStudy (anonymous):

= 3 to the root 15

OpenStudy (anonymous):

@dawnd4fun the key concept is to convert into the same base in order to simplify :)